find the perimeter and the area of the playing area of each table.
50in.by100in,44inby88in.,38in.by76in.

Answer by checkley77(12844)

50in.by100in.
2*50+2*100=100+200=300in. perimeter.
50*100=5,000 in^2 area.
44inby88in.
2*44+2*88=88+176=264in perimeter.
44*88=3,872 in^2 area.
38in.by76in
2*38=2*76=76+152=228in perimeter.
38*76=2,888in^2 area.